# The dimension of sparse random graph orders

### v1

@@ 3 Proof of Theorem 8
We prove the first inequality. Let $U=U[S]\setminus S$, $D=D[S]\setminus S$ and $W=V\setminus(U\cup S\cup D)$. Then $D[S]\cup U[S]=U\cup S\cup D$. It follows that $W$ is exactly the set of elements $y$ such that $y\parallel x$ for every $x\in S$. Let $d_{1}=\dim P[U\cup S\cup D]$ and $d_{2}=\dim P[V\setminus S]$. Let $\{L^{1}_{j}\}_{1\leq j\leq d_{1}}$ be a realiser of $P[U\cup S\cup D]$, and $\{L^{2}_{j}\}_{1\leq j\leq d_{2}}$ be a realiser of $P[V\setminus S]$. Construct linear oderings $\pi$, $\pi^{\prime}$, $\pi^{\prime\prime}$ and $\pi^{\prime\prime\prime}$ on $V$ as follows:
@@
Let =${\mathcal{R}}=\{\pi_{j},\ j\in[d_{1}]\}\cup\{\pi^{\prime}_{j},\ j\in[d_{1}]\}\cup\{\pi^{\prime\prime}_{i},\ i\in[d_{2}]\}\cup\{\pi^{\prime\prime\prime}_{i},\ i\in[d_{2}]\}$. Note that $|{\mathcal{R}}|=2(d_{1}+d_{2})$. Now Theorem 8 follows by Theorem 6 and the following claim.
@@
Proof of Claim 15 There are three cases of $x,y$ such that $x\nleq_{P}y$ in terms of the location of $x$: $x\in S\cup U$; $x\in W$ and $x\in D\setminus U$. We identify $L\in{\mathcal{R}}$ such that $y\leq_{L}U_{P}[x]$ in each case below.
@@
The proof for the second inequality is similar. Let $d_{1}=\dim P[D[S]]$, $d_{2}=\dim P[U[S]]$ and $d_{3}=\dim P[U\cup D]$. Let $\{L^{1}_{j}\}_{1\leq j\leq d_{1}}$ be a realiser of $P[D[S]]$, $\{L^{2}_{j}\}_{1\leq j\leq d_{2}}$ be a realiser for $P[U[S]]$, and $\{L^{3}_{j}\}_{1\leq j\leq d_{3}}$ be a realiser for $P[U\cup D]$. Construct linear orderings $\pi,\pi^{\prime},\pi^{\prime\prime}$ on $U\cup S\cup D$ as follows:
@@
Let =${\mathcal{R}}=\{\pi_{j},\ j\in[d_{2}]\}\cup\{\pi^{\prime}_{i},\ i\in[d_{1}]\}\cup\{\pi^{\prime\prime}_{k},\ k\in[d_{3}]\}$. It suffices to show that Claim 15 holds here as well. For pairs of $(x,y)$ such that $x\nleq_{P}y$ we consider the two cases in terms of location of $x$: $x\in S\cup U$, and $x\in D\setminus(S\cup U)$. We say a permutation $L$ reverses $(x,y)$ if $y<_{L}U[x]$.
@@
4 Dimension of random bipartite POS: proof of Theorem 9
@@ 4.1 Lower bound
(${n\choose\alpha n}^{2}(1-p)^{\alpha^{2}n^{2}}\leq\left(\frac{e}{\alpha}\right)^{2\alpha n}e^{-p\alpha^{2}n^{2}}=\left(\frac{e^{2}}{\alpha^{2}}e^{-c\alpha}\right)^{\alpha n}\leq\exp\left(-c_{n}\epsilon\alpha^{2}_{c_{n}}n\right)=o(1),$
@@
Let $L$ be a linear extension of $P$. Let $\overline{A}$ denote the greatest $\alpha n$ elements in $A$, and let $\underline{B}$ denote the smallest $\alpha n$ elements in $B$ with respect to the total order $L$. By property (b) of Lemma 17, there exists $i\in\overline{A}$ and $j\in\underline{B}$ such that $i<_{P}j$. Let $F$ be the set of incomparable pairs that are reversed by $L$. Suppose $(x,y)\in F$. Then, either $x\geq_{L}i$ or $y\leq_{L}j$, as otherwise the edge $ij$ in $\mathcal{B}(n,n,p)$ would imply that $x\leq_{L}i<_{L}j\leq_{L}y$, contradicting with $(x,y)\in F$. Consequently, for every $(x,y)\in F$, either $x\in\overline{A}$ or $y\in\underline{B}$. It follows then that $|F|\leq|\overline{A}|n+|\underline{B}|n\leq 2\alpha n^{2}$. ∎
@@ 4.2 Upper bound
$\alpha=Kc$ (5)
where $K>0$ is a fixed constant (independent of $c$) that is to be specified later. Let $S_{A}=\{v\in A:\deg(v)\geq\alpha\}$ and $S_{B}=\{v\in B:\deg(v)\geq\alpha\}$, where $\deg(v)$ denotes the degree of $v$ in $\mathcal{B}(n,n,p)$.
@@
Proof of Theorem 9 The lower bounds in parts (a,b) have been demonstrated in Section 4.1. For the upper bound in part (b), by Theorem 7, a.a.s.
@@ 4.3 Proof of Lemma 19
Proof of Claim 21 There are three cases in terms of the degrees of the $i$ vertices in $A$: (a), all of them have degree 2; (b), one of them have degree 3 or 4, and all of the rest have degree 2; (c) two of them have degree 3, and all of the rest have degree 2. It suffices to specify the neighbours of each vertex in $i$ which immediately yields an upper bound $(j^{2})^{i}\leq j^{2+2i}$ for case (a); an upper bound $(j^{2})^{i-1}j^{4}\leq j^{2+2i}$ for case (b), and an upper bound $(j^{2})^{i-2}(j^{3})^{2}\leq j^{2+2i}$ for case (c). As $t=i+j+1$ and $j\geq i-1$, it follows that $t\geq 2i$. Our claimed bound follows immediately.∎
@@
Proof of Lemma 19 Let $L$ be a component of the subgraph $L$ of $\mathcal{B}(n,n,p)$ induced by $U[S_{A}]$. Suppose $L$ has more than one cycle. Then, there is a connected bicyclic subgraph of $L$ satisfying the condition of Lemma 20(b). Hence, we may assume that $L$ is either a tree or is unicyclic. Hence, $\dim P[U[S_{A}]]=\max\{2,\dim P[C]:\text{$C$ is a component of $L$}\}\leq 4$ by Theorems 3 and 4. ∎
@@ 5 Non-bipartite POS: proof of Theorem 10
In this section, $\mathcal{G}_{n}\sim\mathcal{G}(n,p)$, $P\sim P_{\mathcal{G}(n,p=c_{n}/n)}$, and $c>0$ is fixed. We use $<$ to denote the order in natural numbers, and $<_{P}$ to denote the order in $P$. Throughout this section, $q$ denotes $1-p$.
@@
By [5, Lemma 3.1 (iii)], a.a.s. the maximum size of the down-set or up-set of any element in $P$ is $O(\log n)$. This yields part (a). Part (b) follows by a similar proof as in Lemma 17 and notice that
@@
The following lemma is a generalisation of [5, Lemma 3.1] with $V=\{1\}$. The proof is almost identical to [5, Lemma 3.1], and we include it here for completeness.
@@
Given $S\subseteq U$ where $|S|=s$, let $\mathcal{E}_{S}$ denote the event that $U_{P}[V]\cap U=S$. Suppose that $t_{1}<t_{2}<\cdots<t_{s}$ are the elements in $S$ where $<$ denotes the natural order. Let $k_{i}=|\{t\in U:t_{i-1}<t<t_{i}\}|$ for all $2\leq i\leq s$, and $k_{1}=|\{t\in U:t<t_{1}\}|$, $k_{s+1}=|\{t\in U:t>t_{s}\}|$. Then, with ${\boldsymbol{k}}$ denoting $(k_{1},\ldots,k_{s+1})$ and $\|{\boldsymbol{k}}\|$ denoting $\sum_{i=1}^{s+1}k_{i}=|U|-s$,
@@
[…]
@@
[…]
@@
Proof of Theorem 10 Part (a) follows by a similar argument as in the lower bound proof of Theorem 9. By Lemma 22(a), the number of incomparable pairs $(i,j)\in A^{\prime}\times B^{\prime}$ is a.a.s. asymptotic to $\xi^{2}n^{2}$. By Lemma 22(b) and Lemma 24, a.a.s. every linear extension of $P$ can reverse at most $2\beta\xi n^{2}$ pairs of incomparable pairs of elements in $A^{\prime}\times B^{\prime}$. Hence, the size of any realiser must be at least

### v2

@@ 3 Proof of Theorem 8
We prove the first inequality. Let $U=U[S]\setminus S$, $D=D[S]\setminus S$ and $W=V\setminus(U\cup S\cup D)$. Then $D[S]\cup U[S]=U\cup S\cup D$. It follows that $W$ is exactly the set of elements $y$ such that $y\parallel x$ for every $x\in S$. Let $d_{1}=\dim P[U\cup S\cup D]$ and $d_{2}=\dim P[V\setminus S]$. Let $\{L^{1}_{j}\}_{1\leq j\leq d_{1}}$ be a realiser of $P[U\cup S\cup D]$, and $\{L^{2}_{j}\}_{1\leq j\leq d_{2}}$ be a realiser of $P[V\setminus S]$. Construct total orderings $\pi$, $\pi^{\prime}$, $\pi^{\prime\prime}$ and $\pi^{\prime\prime\prime}$ on $V$ as follows:
@@
Let ${\mathcal{R}}=\{\pi_{j},\ j\in[d_{1}]\}\cup\{\pi^{\prime}_{j},\ j\in[d_{1}]\}\cup\{\pi^{\prime\prime}_{i},\ i\in[d_{2}]\}\cup\{\pi^{\prime\prime\prime}_{i},\ i\in[d_{2}]\}$. Note that $|{\mathcal{R}}|=2(d_{1}+d_{2})$. Now Theorem 8 follows by Theorem 6 and the following claim.
@@
Proof of Claim 16. There are three cases of $x,y$ such that $x\nleq_{P}y$ in terms of the location of $x$: $x\in S\cup U$; $x\in W$ and $x\in D\setminus U$. We identify $L\in{\mathcal{R}}$ such that $y\leq_{L}U_{P}[x]$ in each case below.
@@
The proof for the second inequality is similar. Let $d_{1}=\dim P[D[S]]$, $d_{2}=\dim P[U[S]]$ and $d_{3}=\dim P[U\cup D]$. Let $\{L^{1}_{j}\}_{1\leq j\leq d_{1}}$ be a realiser of $P[D[S]]$, $\{L^{2}_{j}\}_{1\leq j\leq d_{2}}$ be a realiser for $P[U[S]]$, and $\{L^{3}_{j}\}_{1\leq j\leq d_{3}}$ be a realiser for $P[U\cup D]$. Construct total orderings $\pi,\pi^{\prime},\pi^{\prime\prime}$ on $U\cup S\cup D$ as follows:
@@
Let ${\mathcal{R}}=\{\pi_{j},\ j\in[d_{2}]\}\cup\{\pi^{\prime}_{i},\ i\in[d_{1}]\}\cup\{\pi^{\prime\prime}_{k},\ k\in[d_{3}]\}$. It suffices to show that Claim 16 holds here as well. For pairs of $(x,y)$ such that $x\nleq_{P}y$ we consider the two cases in terms of location of $x$: $x\in S\cup U$, and $x\in D\setminus(S\cup U)$. We say a permutation $L$ reverses $(x,y)$ if $y<_{L}U[x]$.
@@
4 Dimension of random bipartite POS
4.1 The sparse regime: proof of Theorem 10
@@ 4.1.1 Lower bound
${n\choose\alpha n}^{2}(1-p)^{\alpha^{2}n^{2}}\leq\left(\frac{e}{\alpha}\right)^{2\alpha n}e^{-p\alpha^{2}n^{2}}=\left(\frac{e^{2}}{\alpha^{2}}e^{-c\alpha}\right)^{\alpha n}\leq\exp\left(-c_{n}\epsilon\alpha^{2}_{c_{n}}n\right)=o(1),$
@@
Let $L$ be a linear extension of $P$. Let $\overline{A}$ denote the greatest $\alpha n$ elements in $A$, and let $\underline{B}$ denote the smallest $\alpha n$ elements in $B$ with respect to the total order $L$. By property (b) of Lemma 18, there exist $i\in\overline{A}$ and $j\in\underline{B}$ such that $i<_{P}j$. Let $F$ be the set of incomparable pairs that are reversed by $L$. Suppose $(x,y)\in F$. Then, either $x\geq_{L}i$ or $y\leq_{L}j$, as otherwise the edge $ij$ in $\mathcal{B}(n,n,p)$ would imply that $x\leq_{L}i<_{L}j\leq_{L}y$, contradicting with $(x,y)\in F$. Consequently, for every $(x,y)\in F$, either $x\in\overline{A}$ or $y\in\underline{B}$. It follows then that $|F|\leq|\overline{A}|n+|\underline{B}|n\leq 2\alpha n^{2}$. ∎
@@ 4.1.2 Upper bound
$\alpha=Kc_{n}$ (5)
where $K>0$ is a fixed constant (independent of $c_{n}$) that is to be specified later. Let $S_{A}=\{v\in A:\deg(v)\geq\alpha\}$ and $S_{B}=\{v\in B:\deg(v)\geq\alpha\}$, where $\deg(v)$ denotes the degree of $v$ in $\mathcal{B}(n,n,p)$.
@@
Proof of Theorem 10. The lower bounds in parts (a,b) have been demonstrated in Section 4.1.1. For the upper bound in part (b), by Theorem 7, $a.$s.
@@ 4.1.3 Proof of Lemma 20
Proof of Claim 22. There are three cases in terms of the degrees of the $i$ vertices in $A$: (a), all of them have degree 2; (b), one of them have degree 3 or 4, and all of the rest have degree 2; (c) two of them have degree 3, and all of the rest have degree 2. It suffices to specify the neighbours of each vertex in $i$ which immediately yields an upper bound $(j^{2})^{i}\leq j^{2+2i}$ for case (a); an upper bound $(j^{2})^{i-1}j^{4}\leq j^{2+2i}$ for case (b), and an upper bound $(j^{2})^{i-2}(j^{3})^{2}\leq j^{2+2i}$ for case (c). As $t=i+j+1$ and $j\geq i-1$, it follows that $t\geq 2i$. Our claimed bound follows immediately.∎
@@
Proof of Lemma 20. Let $L$ be a component of the subgraph $L$ of $\mathcal{B}(n,n,p)$ induced by $U[S_{A}]$. Suppose $L$ has more than one cycle. Then, there is a connected bicyclic subgraph of $L$ satisfying the condition of Lemma 21(b). Hence, we may assume that $L$ is either a tree or is unicyclic. Hence, $\dim P[U[S_{A}]]=\max\{2,\dim P[C]:\text{$C$ is a component of $L$}\}\leq 4$ by Theorems 3 and 4. ∎
4.2 The co-sparse regime: proof of Theorem 9
Let $\bar{B}$ denote the bipartite complement of $B\sim\mathcal{B}(n,n,p)$. Then, $\bar{B}\sim\mathcal{B}(n,n,q)$ where q=1pq=1-p.
4.2.1 Proof of part (a)
Claim 23.
Suppose that $q=O(1/n)$ and $q=\omega(1/n^{2})$. A.a.s. $\bar{B}$ contains $\Theta(n^{2}q)$ isolated edges, i.e. components whose order is exactly two.
Proof.
Let $X$ denote the number of isolated edges in $\bar{B}$. Then,
${\mathbb{E}}X=n^{2}q(1-q)^{2(n-1)}=\Theta(n^{2}q)\quad\text{as $q=O(1/n)$},$
and
${\mathbb{E}}X(X-1)=n^{2}(n-1)^{2}q^{2}(1-q)^{4(n-1)-2}\sim({\mathbb{E}}X)^{2}.$
It follows by Chebyshev’s inequality that a.a.s. $X\sim{\mathbb{E}}X$. ∎
We first prove the lower bound on $\dim P_{\mathcal{B}(n,n,p)}$. Let $e_{1},\ldots,e_{m}$ be the set of isolated edges in $\bar{B}$. Then, a.a.s. $m=\Theta(nq^{2})$. Let $(A^{\prime},B^{\prime})\subseteq(A,B)$ be the set of vertices in $\cup_{j=1}^{m}\{e_{j}\}$. Consider the sub-order $P^{\prime}$ of $P$ induced by $A^{\prime}\cup B^{\prime}$. We observe that $P^{\prime}$ is a bipartite order such that every element in $A^{\prime}$ is comparable to all but exactly one element in $B^{\prime}$, and the same holds for every element in $B^{\prime}$. It is well known that the dimension of $P^{\prime}$ is equal to $|A^{\prime}|=m$. Immediately, $\dim P_{\mathcal{B}(n,n,p)}\geq\dim P^{\prime}=\Theta(n^{2}q)$.
Next, we prove the two upper bounds. First we prove that $\dim P_{\mathcal{B}(n,n,p)}=n-\Omega(n)$. Let $A^{\prime}\subseteq A$ be the set of elements that are comparable to every element in $B$, and let $B^{\prime}\subseteq B$ be the set of elements that are comparable to every element in $A$. Let $P^{\prime}$ be the sub-order of $P$ induced by $(A\setminus A^{\prime})\cup(B\setminus B^{\prime})$.
Claim 24.
$\dim P\leq\dim P^{\prime}+1$. Moreover, a.a.s. $|A\setminus A^{\prime}|,|B\setminus B^{\prime}|\sim n(1-p^{n})$.
Proof.
Let $d=\dim P^{\prime}$ and $L_{1},\ldots,L_{d}$ be a realiser of $P^{\prime}$. Let $L$ be an arbitrary total ordering on $A\cup B$, and $L^{\prime}$ be the reverse ordering of $L$. For each $1\leq j\leq d$, extend $L_{j}$ to $A\cup B$ by $L|_{A^{\prime}}<L_{j}|_{(A\setminus A^{\prime})\cup(B\setminus B^{\prime})}<L|_{B^{\prime}}$. Then, add to this set another total ordering defined by $L_{1}|_{(A\setminus A^{\prime})}<L^{\prime}|_{A^{\prime}}<L^{\prime}|_{B^{\prime}}<L_{1}|_{(B\setminus B^{\prime})}$. Obviously this set of $d+1$ total orderings is a realiser for $P$.
Let $X=|A\setminus A^{\prime}|$. Then, ${\mathbb{E}}X=n(1-p^{n})$, and ${\mathbb{E}}X(X-1)=n(n-1)(1-p^{n})^{2}\sim({\mathbb{E}}X)^{2}$. Thus, by Chebyshev’s inequality, a.a.s. $X\sim n(1-p^{n})$. The same statement holds for $|B\setminus B^{\prime}|$ by symmetry. ∎
By the range of $p$, $n(1-p^{n})=n-\Omega(n)$ and thus, $a.$s. $\dim P=n-\Omega(n)$ by Claim 24. Next, we show that a.a.s. $\dim P=O(n^{2}q)$, which holds trivially if $q=\Theta(1/n)$. Suppose that $q=o(1/n)$. Then, by Claim 24, $a.$s. $|A\setminus A^{\prime}|,|B\setminus B^{\prime}|\sim n-ne^{-qn}\sim n^{2}q$, and consequently, $a.$s. $\dim P=O(n^{2}q)$. ∎
4.2.2 Proof of part (b)
Suppose that $q=1-p=\omega(1/n)$ and $q\leq n^{-1+1/7}$. We need the following properties for $\mathcal{B}(n,n,q)$.
Lemma 25.
Let $\epsilon>0$ be fixed. There exists a constant $C>0$ such that a.a.s. $\mathcal{B}(n,n,q)$ has the following properties:
  1. (P1)
    there are $(1+o(1))qn^{2}$ edges;
  2. (P2)
    $d_{v}\leq C(qn+\log n),\quad\text{for all $v\in A\cup B$},$ where $d_{v}$ denotes the degree of $v$ in $\mathcal{B}(n,n,q)$.
  3. (P3)
    there are no subgraphs isomorphic to $K_{2,3}$;
  4. (P4)
    $\sum_{v\in A\cup B}d_{v}{\boldsymbol{1}}_{\{d_{v}\geq(1+\epsilon)qn\}}=o(qn^{2}).$
Let $B\sim{\mathcal{B}(n,n,p)}$ and $\bar{B}$ the bipartite complement of $B$. Then $\bar{B}\sim\mathcal{B}(n,n,q)$. Suppose that $\mathcal{R}$ is a realiser for $P=P_{\mathcal{B}(n,n,p)}$. The following lemma gives a set of properties of $P$ assuming that $\bar{B}$ satisfies properties (P1)–(P4) of Lemma 25.
Lemma 26.
Assume that $\bar{B}$ satisfies Lemma 25 (P1)–(P4). Let $\bar{d}_{v}$ denote the degree of $v$ in $\bar{B}$. Let $L\in\mathcal{R}$. Let $y_{1}$ be the smallest element in $B$ with respect to $L$ and let $x_{k}<_{L}x_{k-1}<_{L}\cdots<_{L}x_{1}$ be the $k$ greatest elements in $A$ with respect to $L$ such that $x_{k}<_{P}y_{1}$ and all elements $x_{k-1},\ldots,x_{1}$ are incomparable to $y_{1}$ in $P$. Then,
  1. (a)
    $k\leq C(qn+\log n)$;
  2. (b)
    If $(x,y)\in A\times B$ is incmparable in $P$ and is reversed by $L$ then $x\in\{x_{1},\ldots,x_{k-1}\}$;
  3. (c)
    For each $1\leq i\leq k-1$, let $R_{i}$ denote the set of $y\in B$ such that $(x_{i},y)$ is incomparable in $P$ and is reversed by $L$. Then, $y||x_{j}$ for all $y\in R_{i}$ and all $x_{j}$ where $j\geq i$;
  4. (d)
    $|\cup_{3\leq j\leq k-1}R_{j}|\leq 1$, $|R_{2}|\leq 2$ if $k\geq 3$, and $|R_{1}|\leq\bar{d}_{x_{1}}$ if $k\geq 2$.
We first complete the proof of part (b) of Theorem 9 assuming Lemmas 25 and 26. Let $\epsilon>0$. We show that a.a.s. any realiser $|\mathcal{R}|$ must satisfy that $|\mathcal{R}|\geq(1-\epsilon)n$. Let $\mathcal{R}=\{L_{1},\ldots,L_{d}\}$ be a realiser where $d=\dim P$. By Lemma 26(b,d) we may assume that if some $L\in\mathcal{R}$ reverses more than $(1+\epsilon/2)qn$ incomparable pairs in $P$ then there is a unique element $x_{L}\in A$ such that
  • (a)
    $L$ reverses at least $(1+\epsilon/2)qn-3$ incomparable pairs $(x_{L},y)$ where $y\in B$.
  • (b)
    $L$ reverses at most 3 other incomparable pairs.
If $L\in\mathcal{R}$ reverses more than $(1+\epsilon/2)qn$ incomparable pairs in $P$, we colour $L\in\mathcal{R}$ red, and colour the corresponding $x_{L}$ red. By Lemma 26(d) and (a), every red element in $A$ must have degree at least $(1+\epsilon/2)qn-3\geq(1+\epsilon/3)qn$ in $\bar{B}$. By Property (P4) of Lemma 25 and Lemma 26(d), the total number of incomparable pairs reversed by the set of red total orderings in $\mathcal{R}$ is $o(qn^{2})$. Hence, by Lemma 26(a), a.a.s. the number of non-red total orderings in $\mathcal{R}$ must be at least
$\frac{(1+o(1))qn^{2}}{(1+\epsilon/2)qn}\geq(1-\epsilon)n,\quad\text{for all sufficiently large $n$},$
which implies that $|\mathcal{R}|\geq(1-\epsilon)n$ as desired. It only remains to prove Lemmas 25 and 26.
Proof of Lemma 25. (P1) and (P2) are standard results in random graph theory following routine concentration arguments, and thus we skip their proofs. Let $X$ denote the number of subgraphs isomorphic to $K_{2,3}$. Then,
${\mathbb{E}}X=2n^{5}q^{6}=o(1),\quad\text{since $q\leq n^{-1+1/7}$}.$
Hence, (P3) holds a.a.s. Finally, noting that $d_{v}$ has the binomial distribution ${\bf Bin}(n,q)$, it is immediate that for every $v\in A\cup B$ and for all sufficiently small $\epsilon>0$,
${\mathbb{E}}\left(d_{v}{\boldsymbol{1}}_{\{d_{v}\geq(1+\epsilon)qn\}}\right)$ $\leq$ $\sum_{j\geq(1+\epsilon)qn}j\binom{n}{j}q^{j}(1-q)^{n-j}$
$\leq$ $\frac{qn}{1-q}e^{-qn}\sum_{j\geq(1+\epsilon)qn-1}\frac{(nq/(1-q))^{j}}{j!}$
$=$ $O\left(qne^{-qn}\left(\frac{nqe}{(1-q)(1+\epsilon)qn}\right)^{(1+\epsilon)qn}\right)$
$=$ $qn\exp\left(-\frac{\epsilon^{2}}{2}qn+O(q+\epsilon^{3})qn\right)=o(qn).$
Therefore,
${\mathbb{E}}\left(\sum_{v\in A\cup B}d_{v}{\boldsymbol{1}}_{\{d_{v}\geq(1+\epsilon)qn\}}\right)=o(qn^{2}),$
and hence (P4) follows by the Markov inequality.∎
Proof of Lemma 26. Part (a) follows from Lemma 25 (P2). For part (b), notice that for any $x\notin\{x_{1},\ldots,x_{k-1}\}$, we must have $x<_{L}x_{k-1}<_{L}y_{1}\leq_{L}y$ and therefore $(x,y)$ cannot be reversed by $L$. For (c), since all elements in $R_{i}$ are smaller than $x_{i}$ in $L$, they are smaller than $x_{j}$ for all $j\leq i$ in $L$, which implies that no elements in $R_{i}$ can be comparable to any of these $x_{j}$. Part (d) now follows from part (c), as otherwise a copy of $K_{2,3}$ must occur in $\bar{B}$, which would contradict with Lemma 25 (P3). ∎
@@ 5 Non-bipartite POS: proof of Theorem 11
In this section, $\mathcal{G}_{n}\sim\mathcal{G}(n,p)$, $P\sim P_{\mathcal{G}(n,p=c_{n}/n)}$, and $c>0$ is fixed. We use << to denote the order in natural numbers, and $_{P}$ to denote the order in $P$. Throughout this section, $q$ denotes 1p1-p.
@@
By [6, Lemma 3.1 (iii)], a.a.s. the maximum size of the down-set or up-set of any element in $P$ is $O(\log n)$. This yields part (a). Part (b) follows by a similar proof as in Lemma 18 and notice that
@@
The following lemma is a generalisation of [6, Lemma 3.1] with $V=\{1\}$. The proof is almost identical to [6, Lemma 3.1], and we include it here for completeness.
@@
Given $S\subseteq U$ where $|S|=s$, let $\mathcal{E}_{S}$ denote the event that $U_{P}[V]\cap U=S$. Suppose that $t_{1}<t_{2}<\cdots<t_{s}$ are the elements in $S$ where << denotes the natural order. Let $k_{i}=|\{t\in U:t_{i-1}<t<t_{i}\}|$ for all $2\leq i\leq s$, and $k_{1}=|\{t\in U:t<t_{1}\}|$, $k_{s+1}=|\{t\in U:t>t_{s}\}|$. Then, with $\boldsymbol{k}$} denoting $(k_{1},\ldots,k_{s+1})$ and $\|{\boldsymbol{k}}\|$ denoting $\sum_{i=1}^{s+1}k_{i}=|U|-s$,
@@
where $Li_{2}(z):=-\int_{0}^{z}y\log(1-y)dy$ is the dilogarithm function, which has the following expansion for all $z\in\mathbb{C}$ where $|z|\leq 1$:
@@
$n^{-1}\log\left(\prod_{j=0}^{tn-1}(1-q^{j+\beta n})\right)$ $=(1+o_{n}(1))\int_{0}^{t}\log(1-e^{-c\beta-cx})dx$
$=(1+o_{n}(1))\int_{e^{-c\beta-ct}}^{e^{-c\beta}}\frac{1}{c}y^{-1}\log(1-y)dy$
@@
Proof of Theorem 11. Part (a) follows by a similar argument as in the lower bound proof of Theorem 10. By Lemma 27(a), the number of incomparable pairs $(i,j)\in A^{\prime}\times B^{\prime}$ is a.a.s. asymptotic to $\xi^{2}n^{2}$. By Lemma 27(b) and Lemma 29, a.a.s. every linear extension of $P$ can reverse at most $2\beta\xi n^{2}$ pairs of incomparable pairs of elements in $A^{\prime}\times B^{\prime}$. Hence, the size of any realiser must be at least